I am having a terrible time trying to figure this out, any help will be greatly appreciated

Bill and Barbara collected the following data in a chemistry experiment when they were in lab. They were comparing two thermometers. The thermometers were under constant conditions except that the heat was changing under each thermometer as time passed. T is the time of the reading. X is the reading of the first thermometer. Y is the reading of the second thermometer. Bill and Barbara are assuming a linear relationship between the readings X and Y on the thermometers and are to find a linear equation relating this (x,y) relationship.
See data collection chart below.
Time starts at t = 0 min.
time
0
10 min.
20 min
30 min.
x
10
15
28
40
y
50
59
82.4
104
a.       Create a graph locating these data points on an (x,y) coordinate plane.
b.      Find the slope of this graph.
c.       Create the linear equation based on the data info collected.

Answer by stanbon(75887)

Bill and Barbara collected the following data in a chemistry experiment when they were in lab. They were comparing two thermometers. The thermometers were under constant conditions except that the heat was changing under each thermometer as time passed. T is the time of the reading. X is the reading of the first thermometer. Y is the reading of the second thermometer. Bill and Barbara are assuming a linear relationship between the readings X and Y on the thermometers and are to find a linear equation relating this (x,y) relationship.
See data collection chart below.
Time starts at t = 0 min.
time
0
10 min.
20 min
30 min.
x
10
15
28
40
y
50
59
82.4
104
a. Create a graph locating these data points on an (x,y) coordinate plane.
b. Find the slope of this graph.
c. Create the linear equation based on the data info collected.
You have more than two points but only need two.
(10,50) (15,59)
-----
slope = (59-50)/(15-10) = 9/5
--------------
Form of you linear equation is y = mx + b
You know y = 50 when x = 10 and you know m = 9/5; solve for "b":
50 = (9/5)*10 + b
b = 50 - 18 = 32
------------------------
Equation:
Y = (9/5)X + 32
===================
Note: This happens to be the formula for converting temperature
in Celsius to temperature in Farenheit.
============================
Cheers,
Stan H.